# Prospects for detecting asteroid-mass primordial black holes in extreme-mass-ratio inspirals with continuous gravitational waves

### v1

@@ I Introduction
The primordial black hole (PBH) hypothesis may explain a variety of puzzling cosmological observations, including the purported existence of DM Hawking (1971); Chapline (1975), LIGO, Virgo and KAGRA binary black hole mergers  Aasi et al. (2015); Acernese et al. (2014); Abbott et al. (2016a, b, c, 2017a, 2017b, 2017c, 2019, 2020a, 2020b, 2020c, 2020d, 2020e), excess microlensing events Alcock et al. (1997), the missing satellite problem Drlica-Wagner et al. (2015), and anisotropies in the cosmic microwave background Bullock and Boylan-Kolchin (2017), without introducing physics beyond the standard model Carr et al. (2021). PBHs could have formed from overdensities in the very early in the universe with masses spanning almost 20 orders of magnitude Villanueva-Domingo et al. (2021):
$M_{\text{PBH}}\sim 10^{6}M_{\odot}\left(\frac{t}{s}\right),$ (1)
@@
If PBHs exist, they could compose a fraction or the totality of dark matter (DM). While the fraction of DM that PBHs could compose, $f_{\text{PBH}}$, has been heavily constrained across a wide parameter space, one particular regime, asteroid-mass ( $10^{-16}M_{\odot}\lesssim M_{\text{PBH}}\lesssim 10^{-12}M_{\odot}$) PBHs, has been able to evade almost all constraints. In this regime, PBHs would be too heavy to avoid evaporating, but too light to amplify the flux from stars that they lens to a high-enough level detectable by microlensing experiments. In particular, if PBHs have masses of $\lesssim 10^{-10}M_{\odot}$, their Schwarzschild radii are approximately equal to, or less than, the wavelength of the light that is being lensed Green and Kavanagh (2021), resulting in a strong suppression of the amplification induced by the lens Ulmer and Goodman (1995) [^1] In principle, the Hyper Suprime-Cam (HSC) experiment Aihara et al. (2018a, b) can probe $f_{\text{PBH}}\lesssim 1$ for lensing masses down to $10^{-12}M_{\odot}$ Niikura et al. (2019), although if the light source has a larger size than assumed in Smyth et al. (2020), the constraints are almost completely relaxed between $\sim[10^{-12},10^{-11}]M_{\odot}$ Smyth et al. (2020)..
@@
One promising avenue could be to search for femtolensing of seconds-long gamma ray bursts (GRBs) Gould (1992). In femtolensing, light rays from a GRB would traverse different path-lengths around the PBH lens and thus result in an interference pattern Davidson and Schwetz (2016). In particular, a lack of observed femtolensing events of GRBs observed by Fermi has led to constraints on $f_{\text{PBH}}\lesssim 0.1$ between $\sim[10^{-16},10^{-14}]M_{\odot}$ Barnacka et al. (2012), although the assumption of point-mass GRB sources has been questioned: if the GRB source has a finite size, photons would be emitted from different locations on the source, which would all follow different paths as they bend around the lens, potentially masking any oscillations visible in the energy spectrum Davidson and Schwetz (2016). Additionally, the large sizes of GRBs relative to the lensing mass in the lensing plane make most unsuitable for such searches, relaxing most projected constraints Katz et al. (2018). Thus, observations of small-size GRBs are needed to produce realistic constraints in the asteroid-mass regime.
@@
Another avenue could be to look at microlensing of light from x-ray pulsars by asteroid-mass PBH lenses, because (1) wave-optics effects are negligible for $M_{\text{PBH}}\gtrsim 10^{-13}M_{\odot}$ Gould (1992), and (2) pulsars are compact objects, which reduces the impact of finite-source size on the lensing of light compared to using extended sources of light Hickox et al. (2004). While it is not currently possible to constrain PBHs using RXTE Levine et al. (1996), future observations with eXTP Zhang et al. (2019) may constrain $f_{\text{PBH}}\lesssim 1$ between $\sim$ [few ×$\times 10^{-16}$, few ×$\times 10^{-13}]M_{\odot}$ Bai and Orlofsky (2019), reaching at best $f_{\text{PBH}}\simeq 0.1$ at $2\times 10^{-15}M_{\odot}$ (see Fig. 5 of Bai and Orlofsky (2019)).
@@
Even if exTP could obtain such constraints in the future, a key mass regime still remains unconstrained that spans almost an order of magnitude, $\sim[10^{-13},10^{-12}]M_{\odot}$, which lies in between the exTP and Subaru/HSC constraints. One possible probe of this mass regime lies in attempting to detect GWs from very slowly inspiraling extreme mass-ratio binaries, consisting of an ordinary compact object with a mass $\mathcal{O}(M_{\odot})$ and a significantly lighter asteroid-mass compact object, i.e. a PBH. Such so-called extreme-mass ratio inspirals (EMRIs) are usually considered in the context of future spaced-based GW detectors, in which an ordinary compact object of $\mathcal{O}(M_{\odot})$ inspirals around and then plunges into a supermassive black hole Babak et al. (2017), though others have considered so-called “mini-EMRI” systems, composed of $\mathcal{O}(1-10)M_{\odot}$ ordinary compact object with a planetary-mass exotic compact object orbiting around it, that would be visible in current ground-based detectors Guo and Miller (2022).
@@
We consider solely the inspiral portion of EMRI systems, specifically at frequencies at which the GW signal appears to be almost monochromatic and ever-lasting. Thus, the signal falls into the category of “continuous waves” (CWs), for which extensive method development has taken place over the last few decades within and outside of the LIGO, Virgo and KAGRA collaborations Sieniawska and Bejger (2019); Tenorio et al. (2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023). However, by only analyzing the inspiral portion of the lifetime of the EMRI system, we pay a price in sensitivity: while searches for merging black holes reach could reach out to at least Gpc, we could, at best, detect systems at the scale of the Galaxy, $\mathcal{O}$(kpc) Miller et al. (2022). Despite the small distance reach of GW searches to such systems, we are still able to project stringent constraints in the asteroid-mass PBH regime.
@@
In this paper, we make a number of considerations regarding the prospects for detection of asteroid-mass PBHs using CW search techniques. While previous works have made specific choices of $m_{1},m_{2}$ and ignored eccentricity in the context of current CW searches Miller et al. (2022); Abbott et al. (2022a), and have not yet been able constrain $f_{\text{PBH}}\leq 1$, we now generalize our thinking to see how constraints on the fraction of DM that PBHs could compose change as a function of (1) GW detector choice, (2) CW search parameters ($T_{\text{FFT}},\dot{f}_{\rm max}$, etc.), (3) source parameters ($m_{1},m_{2}$, eccentricity), and (4) GW frequency bands that are analyzed. Our considerations indicate that searches for CWs can make significant contributions to constraining the asteroid-mass PBH regime, especially if some analysis methods are tuned to consider PBH EMRIs, and even if no changes are made to existing analysis methods.
@@ II Signal model
Two compact objects in orbit around their center of mass will emit GWs as they approach each other. Equating the orbital energy loss with GW power, we can obtain the rate of change of the frequency over time, i.e. the spin-up $\dot{f}$, in the quasi-Newtonian limit (i.e. far from merger) Maggiore (2008):
@@
where $\mathcal{M}\equiv\frac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}$ is the chirp mass of the system, $f_{\text{gw}}$ is the GW frequency, $c$ is the speed of light, and $G$ is Newton’s gravitational constant.
@@
The amplitude $h_{0}(t)$ of the GW signal also evolves with time Maggiore (2008):
@@ III.1 Background
Multiple groups search for CWs emitted from anywhere in the sky Abbott et al. (2022a); Steltner et al. (2023). The parameter space has four dimensions: $f,\dot{f},\alpha,\delta$, where $\alpha,\delta$ refer to the sky position of the source. Despite such a simplistic signal model, these searches cannot be performed fully coherently, since each sky position has to be targeted individually, and the computational cost of such analyses scales with $T_{\text{obs}}^{6}$ Prix (2009). This means that semi-coherent CW search techniques must be employed, which divide the data into smaller chunks of length $T_{\text{FFT}}\ll T_{\text{obs}}$, in order to both save computational costs ($\propto T_{\text{obs}}^{2}T_{\text{FFT}}^{4}$ now) and to increase robustness against theoretical uncertainties in the GW search (e.g. phase coherence could be lost over such long observation times). Computational limitations also require that a fixed range of $\dot{f}$ be searched over, since the number of points in the $\dot{f}$ grid scales with $T_{\text{FFT}}^{4}$ Prix (2009).
When searches for CWs do not find a significant candidate, upper limits on the minimum detectable GW amplitude are produced, usually averaged over the sky but not always Abbott et al. (2022a); Dergachev and Papa (2024), as a function of frequency. These limits are derived either through injecting fake signals and recovering them with a particular method, or through analytic / data-driven procedures that encapsulate the properties of the noise while producing conservative constraints with respect to those that could have been obtained through injections Miller et al. (2021a); Abbott et al. (2022b, a); Miller et al. (2024a, b). An example set of sky-averaged synthetic upper limits that we would obtain on the GW amplitude $h_{0,\rm min}$ using the ET power spectral density is shown Fig. 1, which will be the starting point for our calculations of projected constraints on PBHs.
Figure 1: Synthetic upper limits obtained using Eq. 8 in each 1 Hz band using an ET sensitivity curve and the following analysis parameters: $T_{\text{FFT}}=1.5$ days; $T_{\text{obs}}=1$ year; $\Gamma=0.95$, $CR_{\text{thr}}=5$, and $\theta_{\text{thr}}=2.5$
@@ III.2 Limitations of current searches and constraints on asteroid-mass PBHs
In practice, when performing an all-sky search, preference is usually given to negative values of $\dot{f}$, since the proposed targets of such searches, lumpy isolated neutron stars, should be spinning down, and thus $|\dot{f}|\lesssim 10^{-9}$ Hz/s. The maximum positive $\dot{f}$ searched for limits the extent to which we can use Eq. 4 to approximate Eq. 3. Thus, efforts to constrain PBH abundance cannot typically leverage upper limits at frequencies above $\sim 100$ Hz (see Fig. 1 of Miller et al. (2022)), unless $\mathcal{M}$ is sufficiently small, which, unfortunately, implies smaller $h_{0}$ (Eq. 5). Furthermore, even though in principle PBH abundance for EMRI systems using results from CW searches can be constrained, Refs. Miller et al. (2022); Abbott et al. (2022a) do not yet account for eccentricity of such binaries, and how that would impact the constraints, which we will do in Section VI.
@@ IV Methodology
Following Miller et al. (2022); Miller (2024); Bhattacharya et al. (2024), we compute the sensitivity to inspiraling PBH binaries obtainable with semi-coherent techniques. In our case, we use a formula to compute the sensitivity of the frequency-Hough (a pattern recognition technique used in all-sky searches Astone et al. (2014)) towards quasi-monochromatic sources that are sinusoidal within each $T_{\text{FFT}}$:
@@
×$\times\left(\frac{S_{n}(f_{\text{gw}}=50\text{ Hz})}{3.58\times 10^{-49}\text{ Hz}^{-1}}\right)^{1/2}\left(\frac{\Lambda}{12.81}\right)$ (8)
where $\Lambda=12.81$ for reasonable choices of analysis parameters for the frequency-Hough (see Appendix A).
@@
$d(f)$ $=\frac{1}{\Lambda}\sqrt{\frac{T_{\text{FFT}}^{1/2}T_{\text{obs}}^{1/2}}{{S_{n}(f)}}}\left(\frac{G\mathcal{M}}{c^{2}}\right)^{5/3}\left(\frac{\pi f_{\text{gw}}(t)}{c}\right)^{2/3}$
@@
×$\times\left(\frac{S_{n}(f=50\text{ Hz})}{3.58\times 10^{-49}\text{ Hz}^{-1}}\right)^{-1/2}\left(\frac{\Lambda}{12.81}\right)^{-1}$
×$\times\left(\frac{\mathcal{M}}{10^{-6}M_{\odot}}\right)^{5/3}\left(\frac{f_{\text{gw}}}{50\text{ Hz}}\right)^{2/3}$ (9)
@@
where the brackets indicate the expectation value of the space-time volume. One part of this equation is the Euclidean volume of a sphere, since we do not need to consider any cosmological effects for nearby sources. The other piece, $T$, is: $T=$max$(T_{\text{obs}},\Delta T)$. $\Delta T$ is the time spent by the binary system in a given frequency range $[f,f+\delta f]$:
@@
where $\delta f$ is the spacing in frequency at which we evaluate $d(f)$. Note that we consider this formulation of $T$ because the source durations tend to greatly exceed the observation time, so we are sensitive, in one particular frequency bin, to multiple sources emitting GWs before and/or during $T_{\text{obs}}$ Miller et al. (2022); Bhattacharya et al. (2024). Now, the number of binaries detectable at a given frequency is:
@@
$\mathcal{R}=\frac{3}{4\pi}\left(\sum_{i}\left\langle VT\right\rangle(f_{i})\right)^{-1}.$ (14)
We can equate the rate densities in Eq. 14 to analytic models for formation rate densities of PBHs in the case of asymmetric-mass ratio binaries, with $m_{2}\ll m_{1}$:
@@
×$\times\left(\frac{m_{1}}{M_{\odot}}\right)^{-32/37}\left(\frac{m_{2}}{m_{1}}\right)^{-34/37}\left(f_{\rm PBH}\right)^{53/37}~{},$ (15)
where $f_{\text{sup}}$ is suppression factor induced by the presence of nearby PBHs that could break up binaries, and $f(m_{1})$ and $f(m_{2})$ are the mass functions of the primary and secondary components of the binary, respectively.

### v2

# Prospects for detecting asteroid-mass primordial black holes in extreme mass-ratio inspirals with continuous gravitational waves

@@ I Introduction
The primordial black hole (PBH) hypothesis may explain a variety of puzzling cosmological observations, including the purported existence of DM Hawking:1971ei; Chapline:1975ojl, LIGO, Virgo and KAGRA binary black hole mergers  aasi2015advanced; acernese2014advanced; Abbott:2016blz; TheLIGOScientific:2016pea; Abbott:2016nmj; Abbott:2017vtc; Abbott:2017oio; Abbott:2017gyy; LIGOScientific:2018mvr; LIGOScientific:2020stg; Abbott:2020uma; Abbott:2020khf; Abbott:2020tfl; Abbott:2020mjq, excess microlensing events MACHO:1996qam, the missing satellite problem DES:2015zwj, and anisotropies in the cosmic microwave background Bullock:2017xww, without introducing physics beyond the standard model Carr:2019kxo. PBHs could have formed from overdensities in the very early in the universe Khlopov:2008qy with masses given by Villanueva-Domingo:2021spv:
$M_{\text{PBH}}\sim 5\times 10^{4}M_{\odot}\left(\frac{t}{1\text{ s}}\right),$ (1)
@@
If PBHs exist, they could compose a fraction or the totality of dark matter (DM) and particular observation signatures Belotsky:2014kca; Clesse:2017bsw; Belotsky:2018wph. While the fraction of DM that PBHs could compose, $f_{\text{PBH}}$, has been heavily constrained across a wide parameter space, one particular regime, asteroid-mass ( $10^{-16}M_{\odot}\lesssim M_{\text{PBH}}\lesssim 10^{-12}M_{\odot}$) PBHs, has been able to evade almost all constraints. In this regime, PBHs would be too heavy to avoid evaporating [^1] AMEGO, however, may be able to detect evaporating asteroid-mass PBHs emitting MeV photons Ray:2021mxu. , but too light to amplify the flux from stars that they lens to a high-enough level detectable by microlensing experiments. In particular, if PBHs have masses of $\lesssim 10^{-10}M_{\odot}$, their Schwarzschild radii are approximately equal to, or less than, the wavelength of the light that is being lensed Green:2020jor, resulting in a strong suppression of the amplification induced by the lens Ulmer:1994ij [^2] In principle, the Hyper Suprime-Cam (HSC) experiment Aihara:2017paw; Aihara:2017tri can probe $f_{\text{PBH}}\lesssim 1$ for lensing masses down to $10^{-12}M_{\odot}$ Niikura:2017zjd, although if the light source has a larger size than assumed in Smyth:2019whb, the constraints are almost completely relaxed between $\sim[10^{-12},10^{-11}]M_{\odot}$ Smyth:2019whb..
@@
One promising avenue could be to search for femtolensing of seconds-long gamma ray bursts (GRBs) gould1992femtolensing. In femtolensing, light rays from a GRB would traverse different path-lengths around the PBH lens and thus result in an interference pattern Davidson:2016uok. In particular, a lack of observed femtolensing events of GRBs observed by Fermi has led to constraints on $f_{\text{PBH}}\lesssim 0.1$ between $\sim[10^{-16},10^{-14}]M_{\odot}$ Barnacka:2012bm, although the assumption of point-mass GRB sources has been questioned: if the GRB source has a finite size, photons would be emitted from different locations on the source, which would all follow different paths as they bend around the lens, potentially masking any oscillations visible in the energy spectrum Davidson:2016uok. Additionally, the large sizes of GRBs relative to the lensing mass in the lensing plane make most unsuitable for such searches, relaxing most projected constraints Katz:2018zrn. Thus, observations of small-size GRBs are needed to produce realistic constraints in the asteroid-mass regime.
@@
Another avenue could be to look at microlensing of light from x-ray pulsars by asteroid-mass PBH lenses, because (1) wave-optics effects are negligible for $M_{\text{PBH}}\gtrsim 10^{-13}M_{\odot}$ gould1992femtolensing, and (2) pulsars are compact objects, which reduces the impact of finite-source size on the lensing of light compared to using extended sources of light Hickox:2004fy. While it is not currently possible to constrain PBHs using RXTE Levine:1996du, future observations with eXTP eXTP:2018anb may constrain $f_{\text{PBH}}\lesssim 1$ between $\sim$ [few $\times 10^{-16}$, few $\times 10^{-13}]M_{\odot}$ Bai:2018bej, reaching at best $f_{\text{PBH}}\simeq 0.1$ at $2\times 10^{-15}M_{\odot}$ (see Fig. 5 of Bai:2018bej).
Even if eXTP could obtain such constraints in the future, a key mass regime still remains unconstrained that spans almost an order of magnitude, $\sim[10^{-13},10^{-12}]M_{\odot}$, which lies in between the eXTP and Subaru/HSC constraints.
@@
One possible probe of this mass regime lies in attempting to detect GWs from very slowly inspiraling extreme mass-ratio binaries, consisting of an ordinary compact object with a mass $\mathcal{O}(M_{\odot})$ and a significantly lighter asteroid-mass compact object, i.e. a PBH. Such so-called extreme-mass ratio inspirals (EMRIs) are usually considered in the context of future spaced-based GW detectors, in which an ordinary compact object of $\mathcal{O}(M_{\odot})$ inspirals around and then plunges into a supermassive black hole Babak:2017tow, though others have considered so-called “mini-EMRI” systems, composed of $\mathcal{O}(1-10)M_{\odot}$ ordinary compact object with a planetary-mass exotic compact object orbiting around it, that would be visible in current ground-based detectors Guo:2022sdd.
Additionally, asteroid-mass PBHs could be constrained via their imprints on a stochastic GW background. If PBHs are generated from overdensities in the early universe, they will lead to scalar-induced GWs, which, if not detected by future space-based detectors, such as LISA, would also stringently constrain asteroid-mass PBHs Yuan:2019udt; Meng:2022low.
@@
Here, consider solely the inspiral portion of EMRI systems, specifically at frequencies at which the GW signal appears to be almost monochromatic and ever-lasting. Thus, the signal falls into the category of “continuous waves” (CWs), for which extensive method development has taken place over the last few decades within and outside of the LIGO, Virgo and KAGRA collaborations Sieniawska:2019hmd; Tenorio:2021wmz; Riles:2022wwz; Piccinni:2022vsd; Miller:2023qyw; Wette:2023dom. However, by only analyzing the inspiral portion of the lifetime of the EMRI system, we pay a price in sensitivity: while searches for merging black holes reach could reach out to at least Gpc, we could, at best, detect systems at the scale of the Galaxy, $\mathcal{O}$(kpc) Miller:2021knj. Despite the small distance reach of GW searches to such systems, we are still able to project stringent constraints in the asteroid-mass PBH regime.
@@
In this paper, we make a number of considerations regarding the prospects for detection of asteroid-mass PBHs using CW search techniques. While previous works have made specific choices of $m_{1},m_{2}$ and ignored eccentricity in the context of current CW searches Miller:2021knj; KAGRA:2022dwb, and have not yet been able constrain $f_{\text{PBH}}\leq 1$, we now generalize our thinking to see how constraints on the fraction of DM that PBHs could compose change as a function of (1) GW detector choice, (2) CW search parameters ($T_{\text{FFT}},\dot{f}_{\rm max}$, etc.), (3) source parameters ($m_{1},m_{2}$, eccentricity), and (4) GW frequency bands that are analyzed. Our considerations indicate that searches for CWs can make significant contributions to constraining the asteroid-mass PBH regime, especially if some analysis methods are tuned to consider PBH EMRIs, and even if no changes are made to existing analysis methods.
@@ II Signal model
Two compact objects in orbit around their center of mass will emit GWs as they approach each other. Equating the orbital energy loss with GW power, we can obtain the rate of change of the frequency over time, i.e. the spin-up $\dot{f}$, in the quasi-Newtonian limit (i.e. far from merger) maggiore2008gravitational:
@@
where $\mathcal{M}\equiv\frac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}$ is the chirp mass of the system, $f_{\text{gw}}$ is the GW frequency, $c$ is the speed of light, and $G$ is Newton’s gravitational constant. Throughout the paper, unless we explicitly say we are considering eccentricity, the results assume quasi-circular orbits.
@@
The amplitude $h_{0}(t)$ of the GW signal also evolves with time maggiore2008gravitational:
@@ III.1 Background
Multiple groups search for CWs emitted from anywhere in the sky KAGRA:2022dwb; Steltner:2023cfk. The parameter space for all-sky searches for isolated neutron stars typically has four dimensions: $f,\dot{f},\alpha,\delta$, where $\alpha,\delta$ refer to the sky position of the source. Despite such a simplistic signal model, these searches cannot be performed fully coherently, since each sky position has to be targeted individually, and the computational cost of such analyses scales with $T_{\text{obs}}^{6}$ Prix:2009oha. This means that semi-coherent CW search techniques must be employed, which divide the data into smaller chunks of length $T_{\text{FFT}}\ll T_{\text{obs}}$, in order to both save computational costs ($\propto T_{\text{obs}}^{2}T_{\text{FFT}}^{4}$ now) and to increase robustness against theoretical uncertainties in the GW search (e.g. phase coherence could be lost over such long observation times). Computational limitations also require that a fixed range of $\dot{f}$ be searched over, since the number of steps in the $\dot{f}$ grid scales with $T_{\text{FFT}}T_{\text{obs}}$Prix:2009oha.
When searches for CWs do not find a significant candidate, upper limits on the minimum detectable GW amplitude are produced, usually averaged over the sky but not always KAGRA:2022dwb; Dergachev:2024knd, as a function of frequency. These limits are derived either through injecting fake signals and recovering them with a particular method, or through analytic / data-driven procedures that encapsulate the properties of the noise while producing conservative constraints with respect to those that could have been obtained through injections Miller:2020vsl; LIGOScientific:2021odm; KAGRA:2022dwb; Miller:2024fpo; Miller:2024jpo. An example set of sky-averaged synthetic upper limits that we would obtain on the GW amplitude $h_{0,\rm min}$ using the ET power spectral density is shown Fig. 1, which will be the starting point for our calculations of projected constraints on PBHs.
Figure 1: Synthetic upper limits obtained using Eq. 8 in each 1 Hz band using an ET sensitivity curve and the following analysis parameters: $T_{\text{FFT}}=1.5$ days; $T_{\text{obs}}=1$ year; $\Gamma=0.95$, $CR_{\text{thr}}=5$, and $\theta_{\text{thr}}=2.5$ .
@@ III.2 Limitations of current searches and constraints on asteroid-mass PBHs
In practice, when performing an all-sky search, preference is usually given to negative values of $\dot{f}$, since the proposed targets of such searches, lumpy isolated neutron stars, should be spinning down, and thus $|\dot{f}|\lesssim 10^{-9}$ Hz/s. The maximum positive $\dot{f}$ searched for limits the extent to which we can use Eq. 4 to approximate Eq. 3. Thus, efforts to constrain PBH abundance cannot typically leverage upper limits at frequencies above $\sim 100$ Hz (see Fig. 1 of Miller:2021knj), unless $\mathcal{M}$ is sufficiently small, which, unfortunately, implies smaller $h_{0}$ (Eq. 5). Furthermore, even though in principle PBH abundance for EMRI systems using results from CW searches can be constrained, Refs. Miller:2021knj; KAGRA:2022dwb do not yet account for eccentricity of such binaries, and how that would impact the constraints, which we will do in Section VI.
@@ IV Methodology
Following Miller:2021knj; Miller:2024rca; Bhattacharya:2024pmp, we compute the sensitivity to inspiraling PBH binaries obtainable with semi-coherent techniques. In our case, we use a formula to compute the sensitivity of the frequency-Hough (a pattern recognition technique used in all-sky searches Astone:2014esa) towards quasi-monochromatic sources that are sinusoidal within each $T_{\text{FFT}}$:
@@
$\times\left(\frac{S_{n}(f_{\text{gw}}=50\text{ Hz})}{3.58\times 10^{-49}\text{ Hz}^{-1}}\right)^{1/2}\left(\frac{\Lambda}{12.81}\right)$ (8)
where $\Lambda=12.81$ based on the official sensitivity of the frequency-Hough, see Eq. 67 of Astone:2014esa. Note that $T_{\text{FFT}}=1.5$ days is an assumption: we do not know what computational resources we will have in the next-generation detector era, but as in Branchesi:2023mws, we assume that we will be able to take longer ones than what we use now (e.g. $T_{\text{FFT}}=8192$ s at frequencies below 128 Hz Astone:2014esa). Using $T_{\text{FFT}}=8192$ s would result in a factor of 2 worse sensitivity in $h_{\text{0,min}}$, thus a factor of 8 smaller on the rate density projected constraints presented below.
@@
$d(f_{\text{gw}})$ $=\frac{1}{\Lambda}\sqrt{\frac{T_{\text{FFT}}^{1/2}T_{\text{obs}}^{1/2}}{{S_{n}(f)}}}\left(\frac{G\mathcal{M}}{c^{2}}\right)^{5/3}\left(\frac{\pi f_{\text{gw}}}{c}\right)^{2/3}$
@@
$\times\left(\frac{S_{n}(f_{\text{gw}}=50\text{ Hz})}{3.58\times 10^{-49}\text{ Hz}^{-1}}\right)^{-1/2}\left(\frac{\Lambda}{12.81}\right)^{-1}$
$\times\left(\frac{\mathcal{M}}{10^{-6}M_{\odot}}\right)^{5/3}\left(\frac{f_{\text{gw}}}{50\text{ Hz}}\right)^{2/3}$ (9)
It is also worth quoting these quantities in method-independent parameters, particularly the sensitivity depth $\mathcal{D}=\sqrt{S_{n}}/h_{0,\rm min}$ that would allow interested readers to, with their own methods, assess their sensitivity to asteroid-mass PBHs. In terms of $\mathcal{D}$:
$d(f_{\text{gw}})=$ $30\,\text{pc}\,\left(\frac{\mathcal{M}}{10^{-6}\,M_{\odot}}\right)^{5/3}\left(\frac{f_{\text{gw}}}{50\,\text{Hz}}\right)^{2/3}$
$\times$ $\left(\frac{\mathcal{D}}{111\,\text{Hz}^{-1/2}}\right)\left(\frac{6\times 10^{-25}\text{ Hz}^{-1/2}}{\sqrt{S_{n}}}\right)$ (10)
@@
[…]
@@
where $\delta f$ is the spacing in frequency at which we evaluate $d(f)$. Note that we consider this formulation of $T$ because the source durations tend to greatly exceed the observation time, so we are sensitive, in one particular frequency bin, to multiple sources emitting GWs before and/or during $T_{\text{obs}}$ Miller:2021knj; Bhattacharya:2024pmp. $\delta f$ is essentially the spacing in frequency at which the upper limits is computed in a real search, when such upper limits are computed with injections. It is not an infinitesimally small quantity, and determines how “monochromatic” the upper limit at each frequency is. Physically, $\delta f$ contains overlapping PBH inspirals that are all at the “same” frequency (within $\delta f$). As we show in Appendix A, it does matter how big or small this quantity is, because in the regime in which the source durations greatly exceed the observation time, we are sensitive, in $\delta f$, to multiple sources emitting GWs.
Now, the number of binaries detectable at a given frequency is:
@@
$\mathcal{R}=\left(\sum_{i}\left\langle VT\right\rangle(f_{i})\right)^{-1}.$ (15)
We can equate the rate densities in Eq. 15 to analytic models Clesse:2015wea; Clesse:2016vqa for formation rate densities of PBHs in the case of asymmetric-mass ratio binaries, with $m_{2}\ll m_{1}$:
@@
$\times\left(\frac{m_{1}}{M_{\odot}}\right)^{-32/37}\left(\frac{m_{2}}{m_{1}}\right)^{-34/37}\left(f_{\rm PBH}\right)^{53/37}~,$ (16)
where $f_{\text{sup}}$ is suppression factor induced by the presence of nearby PBHs that could break up binaries, $f(m_{1})$ and $f(m_{2})$ are the mass functions of the primary and secondary components of the binary, respectively, and $f_{\text{PBH}}$ is the fraction of DM that PBHs could compose.